write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. The solutions are -4,2
Answer by mananth(16946)

(x+4)(x-2)=0
x^2-2x+4x-8=0
x^2+2x-8=0